OpenStudy (anonymous):

use double angle identities 2 sin75^0 cos 75^0

OpenStudy (anonymous):

I have sin 2x = sin2(75) cos(75)

OpenStudy (anonymous):

i don't know how to edit

OpenStudy (anonymous):

i have sin 150 but i don't know what to do with cos75

OpenStudy (anonymous):

hey, sin 2x = 2sinx cos x

OpenStudy (anonymous):

what did i do wrong?

OpenStudy (anonymous):

2(75) cos(75)

OpenStudy (anonymous):

you miss 2 before sin x cos x

OpenStudy (anonymous):

ok, Madina, You wrong at the beginning , You have the form already , you have 2 sin 75 cos 75 = sin 2*75= sin 150 . look up the value of sin 150 plug there and you got the answer

OpenStudy (anonymous):

\[-\sqrt{3}/2\]

OpenStudy (anonymous):

cannot be that, 150 at quadrant 2, sin in that region cannot be negative

OpenStudy (anonymous):

yes cos is negative there

OpenStudy (anonymous):

but yours is end up at sin

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

you calculate for sin 150 , it cannot be - sqr .....

OpenStudy (anonymous):

i honestly don't see what i am missing

OpenStudy (anonymous):

you don't have to do anything, why you calculate cos 75?

OpenStudy (anonymous):

if it is sin, then 1/2

OpenStudy (anonymous):